# FALCON: Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN

### v2

@@ 4.3 Long-Term Memory Feedback
4.4 Short-Term Memory Feedback
@@
where $R_{\text{fine}}(\ast)$ represents the reward coefficient, and $S_{\text{fine}}$ and $E_{\text{fine}}$ denote the start and end positions of the code snippet, respectively. These values are determined based on different types of feedback. Compiler Feedback. For compiler feedback, we adopt the same settings as CodeRL:
@@
$S_{\text{coarse}}=0,\quad E_{\text{coarse}}=T$
where $R_{\text{coarse}}$ is based on compiler feedback with the start and end positions set to 0 and $T$.
@@
Coding Style Feedback. To further enhance the quality of the generated code, we employ AI Feedback to optimize coding style. An evaluation model scores the generated code based on adherence to the expected coding style standards. The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code. The coding style assessment template is provided in Table 8.
Complexity Feedback. Just like with coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores. The complexity assessment template is provided in Table 9.
Long-term Error Type Feedback. Introducing a reward mechanism that combines short-term memory recall error rate, current test error, and long-term memory recall of past task performance enables the model to dynamically adjust its rewards based on past error patterns, adapt to various error types and feedback, and ultimately enhance its generalization ability:
@@
where $N_{\text{error}}$ represents the short-term memory recall, the frequency of each error type in the generated code by the model. This is the immediate feedback for the current task. $P_{\text{error}}$ represents the long-term memory recall, the proportion of each error type. This is the model’s performance statistics on long-term tasks, providing a cumulative history of various error types. By correct rewarding, the model can reduce the occurrence of these errors and thereby enhance the accuracy and quality of the generated code.
5 Experiment
5.1 Quantitative Evaluation on APPS
To ensure a fair comparison, we use the CodeT5 770M model as our baseline. Our benchmarks include the latest advancements that integrate reinforcement learning (RL) with large language models (LLMs), particularly CodeRL, PPOCoder, and RLTF. For evaluation, we apply the same benchmarks and settings used in these previous works. As shown in Table 1 for the experimental results, our FALCON approach delivers additional performance improvements and surpasses other RL-based methods, indicating that RL with appropriate feedback can effectively improve the model output space and thereby enhance the quality of code generation. In particular, our method achieves the highest pass@1 rates of 8.60%, 2.56%, and 1.25% in the Introductory, Interview, and Competition categories, respectively.
Table 1: Quantitative evaluation on the APPS benchmark. “Intro”: introductory, “Inter”: interview, “Comp”: competition-level tasks. To ensure a fair comparison, we apply these RL-based methods, including PPOCoder, CodeRL, and RLTF, using the same base model, CodeT5, as a backbone. We also compare with models that have a larger number of parameters.
@@ 5.1 Quantitative Evaluation on APPS
5.2 Quantitative Evaluation on HumanEval and MBPP
To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model, trained with our method on our custom dataset, using the well-established MBPP and HumanEval benchmarks. We also compare these results against other reinforcement learning methods, such as PPOCoder and RLTF. The experimental results are illustrated in Table 2.
Table 2: The results of pass@1 on the MBPP and HumanEval benchmarks.
@@ 5.2 Quantitative Evaluation on HumanEval and MBPP
5.3 Quantitative Evaluation on CODAL-Bench
Figure 3: Quantitative evaluation on CODAL-Bench
5.4 Quantitative Evaluation on SciCode
Table 3: The results of pass@1 on SciCode benchmarks with and without FALCON.
Model Size Method Subproblem Main Problem
CodeLlama 70B - 10.4 0
CodeLlama 7B w/o 0.4 0
CodeLlama 7B w 3.5 0
DeepSeek-Coder 6.7B w/o 5.2 0
DeepSeek-Coder 6.7B w 8.3 0
5.5 Quantitative Evaluation on Agentbench
To further evaluate the performance of our framework, we conduct a systematic assessment on AgentBench Liu et al. (2023b), focusing specifically on long-term memory capabilities. Since our primary focus is on code generation tasks, we select three environments within AgentBench: Operating System (OS), Database (DB), and Knowledge Graph (KG). In this evaluation, we compare proprietary models (such as GPT-4 and GPT-3.5) with open-source models (such as Codellama and Qwen 2.5) Liang et al. (2024). The results reveal that the models optimized through our framework exhibit significant improvements, particularly in the OS environment with an increase of 4.8 percentage points. The experimental results are illustrated in Table 4.
Table 4: Test set results of AGENTBENCH.
@@ 5.5 Quantitative Evaluation on Agentbench
5.6 Ablation Studies
The Influence of Models. To validate the scalability and robustness of our framework, we conduct experiments with the larger model, DeepSeek-Coder-Instruct-6.7B, to further evaluate its performance. Notably, the improvements in introductory-level tasks are significant, which can be attributed to the use of long-term memory that enhances the quality of generated data and further unlocks the model’s potential. The results are illustrated in Table 5.
Table 5: Different large language models as the backbone
@@ 5.6 Ablation Studies
The Influence of Memory. To validate the impact of long-term and short-term memories on code generation capabilities, we conduct ablation experiments using CodeT5 as the base model and test it on the APPS dataset. As shown in Table 6, the experimental results indicate that both long- and short-term memory feedback enhances the model’s code generation performance effectively, while the short-term memory feedback demonstrates a more significant improvement. This improvement can be attributed to the effective reward design which plays a positive role in fine-tuning the model.
Table 6: Effect of long and short memories on different performance metrics
@@
6 Conclusions and Future Work
In this work, we propose FALCON, a novel RL-based feedback He et al. (2024) framework that enhances automated code generation by integrating long-term and short-term memory feedbacks within a meta-reinforcement learning strategy to better capture human intent like RL based preference learning. Long-term memory retains past interactions to improve code quality and reduce repetitive mistakes, while short-term memory enables immediate adjustments based on recent feedback from compilers and AI systems. This dual-memory approach addresses limitations in existing models that struggle to align code generation with user intent, especially in specialized tasks or edge cases. By incorporating non-differentiable code features like style and complexity into the feedback loop, FALCON ensures that the generated code is not only functionally correct but also adheres to real-world programming standards. Extensive evaluations on benchmarks including APPS, HumanEval, and CODAL-Bench demonstrate that FALCON outperforms existing RL-based methods, achieving higher functional correctness and better coding style adherence. In future work, we aim to expand FALCON’s capabilities by incorporating a broader diversity of programming languages and tackling more complex code generation challenges.
@@ Appendix C Prompts
Table 7: Code generation template
@@
Table 8: Coding style assessment template.
Coding Style Assessment
Reward Scale: Rate outputs on a scale of -1 to 2
Table 9: Complexity assessment template.
Complexity Assessment
Reward Scale: Rate outputs on a scale of -1 to 2
Table 10: Instruction following assessment template.
Instruction Following Assessment
Evaluation Criteria
Reward Scale: Rate outputs on a scale of -1 to 2
Appendix D Error Category
Due to the differences in languages accepted by Compiler Feedback during unit tests for various language tasks, we have standardized the definition of sub-errors in Compiler Feedback. The table 11 12 13 below outlines our specifications for Python, C, and Java.
Table 11: Common Python errors with categories.
Sub-error Description Category
Indentation Error Wrong indentation format. Syntax Error
Index Error Index operation is out of bounds. Runtime Error
Timeout Error Code execution time exceeds time limit. Runtime Error
Name Error A local or global name is not defined. Runtime Error
Import Error The imported package is not found. Runtime Error
Table 12: Common C Language errors with categories.
Sub-error Description Category
Division by Zero Attempting to divide a number by zero. Runtime Error
Table 13: Common Java Language errors with categories.
Sub-error Description Category
Appendix E Supplementary Experiments
E.1 Quantitative Evaluation on APPS
Table 14: Results of pass@1 on APPS.
@@ E.1 Quantitative Evaluation on APPS
E.2 Quantitative Evaluation on LeetCode
Table 15: Results of pass@1 on LeetCode.

### v3

@@ IV-C Long-Term Memory Feedback
IV-D Short-Term Memory Feedback
@@
where $R_{\text{fine}}(\ast)$ represents the reward coefficient, and $S_{\text{fine}}$ and $E_{\text{fine}}$ denote the start and end positions of the code snippet, respectively. These values are determined based on different types of feedback.
Compiler Feedback. For compiler feedback, we adopt the same settings as CodeRL:
@@
where $R_{\text{coarse}}$ is based on compiler feedback with the start position $S_{\text{coarse}}=0$ and end position $E_{\text{coarse}}=T$.
@@
Coding Style Feedback. To further enhance the quality of the generated code, we employ AI Feedback to optimize coding style. An evaluation model scores the generated code based on adherence to the expected coding style standards. The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code. The coding style assessment template is provided in Table X.
Complexity Feedback. Similar to the coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores. The complexity assessment template is provided in Table XI.
Long-term Error Type Feedback. Introducing a reward mechanism that combines short-term memory recall error rate, current test error, and long-term memory feedback of past task performance enables the model to dynamically adjust its rewards based on past error patterns, adapt to various error types and feedback, and ultimately enhance its generalization ability:
@@
where $N_{\text{error}}$ represents the frequency of each error type in the generated code, which is the short-term memory feedback and the immediate feedback for the current task. $P_{\text{error}}$ represents the proportion of each error type, which is the long-term memory feedback. This is the model’s performance statistics on long-term tasks, providing a cumulative history of various error types. By correct rewarding, the model can reduce the occurrence of these errors and thereby enhance the accuracy and quality of the generated code.
TABLE I: Quantitative evaluation on the APPS benchmark. “Intro”: introductory, “Inter”: interview, “Comp”: competition-level tasks. To ensure a fair comparison, we apply these RL-based methods, including PPOCoder, CodeRL, and RLTF, using the same base model, CodeT5, as a backbone. We also compare with models that have a larger number of parameters.
@@
V Experiment
V-A Quantitative Evaluation on APPS
To ensure a fair comparison, we use the CodeT5 770M model as our baseline. Our benchmarks include the latest advancements that integrate reinforcement learning (RL) with large language models (LLMs), particularly CodeRL, PPOCoder, and RLTF. For evaluation, we apply the same benchmarks and settings used in these previous works. As shown in Table I for the experimental results, our FALCON approach delivers additional performance improvements and surpasses other RL-based methods, indicating that RL with appropriate feedback can effectively improve the model output space and thereby enhance the quality of code generation. In particular, our method achieves the highest pass@1 rates of 8.60%, 2.56%, and 1.25% in the Introductory, Interview, and Competition categories, respectively.
V-B Quantitative Evaluation on HumanEval and MBPP
To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model trained with our method on APPS, using the well-established MBPP and HumanEval benchmarks. We also compare these results against other reinforcement learning methods, such as PPOCoder and RLTF. The experimental results are illustrated in Table II.
TABLE II: The results of pass@1 on the MBPP and HumanEval benchmarks.
@@ V-B Quantitative Evaluation on HumanEval and MBPP
TABLE III: Test set results of AGENTBENCH.
@@
V-C Quantitative Evaluation on CODAL-Bench
Figure 3: Quantitative evaluation on CODAL-Bench
V-D Quantitative Evaluation on SciCode
TABLE IV: The results of pass@1 on SciCode benchmarks with and without FALCON.
Model Size Method Subproblem Main Problem
CodeLlama 70B - 10.4 0
CodeLlama 7B w/o 0.4 0
CodeLlama 7B w 3.5 0
DeepSeek-Coder 6.7B w/o 5.2 0
DeepSeek-Coder 6.7B w 8.3 0
V-E Quantitative Evaluation on Agentbench
To further evaluate the performance of our framework, we conduct a systematic assessment on AgentBench [23], focusing specifically on long-term memory capabilities. Since our primary focus is on code generation tasks, we select three environments within AgentBench: Operating System (OS), Database (DB), and Knowledge Graph (KG). In this evaluation, we compare proprietary models (such as GPT-4 and GPT-3.5) with open-source models (such as Codellama and Qwen 2.5) [24]. The results reveal that the models optimized through our framework exhibit significant improvements, particularly in the OS environment with an increase of 4.8 percentage points. The experimental results are illustrated in Table III.
V-F Ablation Studies
The Influence of Models. To validate the scalability and robustness of our framework, we conduct experiments with DeepSeek-Coder-Instruct-6.7B. Notably, the improvements in introductory-level tasks are significant, which can be attributed to the use of long-term memory that enhances the quality of generated data and further unlocks the model’s potential. The results are illustrated in Table V.
TABLE V: Different large language models as the backbone
@@ V-F Ablation Studies
The Influence of Memory. To validate the impact of long-term and short-term memories on code generation capabilities, we conduct ablation experiments using CodeT5 as the base model and test it on the APPS dataset. As shown in Table VI, the experimental results indicate that both long- and short-term memories enhance the model’s code generation performance effectively, while the short-term memory feedback demonstrates a more significant improvement attributed to the reward design which plays an important role in fine-tuning the model.
TABLE VI: Effect of long and short memories on different performance metrics
@@
VI Conclusions and Future Work
In this work, we propose FALCON, a novel RL-based feedback framework that enhances automatic code generation by integrating long-term and short-term memory feedback within a meta-reinforcement learning strategy to better capture human intent like RL based preference learning. Long-term memory retains past interactions to improve code quality and reduce repetitive mistakes, while short-term memory enables immediate adjustments based on recent feedback from compilers and AI systems. This dual-memory approach addresses limitations in existing models that struggle to align code generation with user intent, especially in specialized tasks or edge cases. By incorporating non-differentiable code features like style and complexity into the feedback loop, FALCON ensures that the generated code is not only functionally correct but also adheres to real-world programming standards. Extensive evaluations on benchmarks including APPS, HumanEval, and CODAL-Bench demonstrate that FALCON outperforms existing RL-based methods, achieving higher functional correctness and better coding style adherence. In future work, we aim to expand FALCON’s capabilities by incorporating a broader diversity of programming languages and tackling more complex code generation challenges.
@@ Appendix C Prompts
TABLE VII: Code generation template
@@
TABLE VIII: Results of pass@1 on APPS.
@@
TABLE IX: Results of pass@1 on LeetCode.
@@
TABLE X: Coding style assessment template.
Coding Style Assessment
Reward Scale: Rate outputs on a scale of -1 to 2
TABLE XI: Complexity assessment template.
Complexity Assessment
Reward Scale: Rate outputs on a scale of -1 to 2
TABLE XII: Instruction following assessment template.
Instruction Following Assessment
Evaluation Criteria
Reward Scale: Rate outputs on a scale of -1 to 2
Appendix D Error Category
Due to the differences in languages accepted by Compiler Feedback during unit tests for various language tasks, we have standardized the definition of sub-errors in Compiler Feedback. The table XIII XIV XV below outlines our specifications for Python, C, and Java.
TABLE XIII: Common Python errors with categories.
Sub-error Description Category
Indentation Error Wrong indentation format. Syntax Error
Index Error Index operation is out of bounds. Runtime Error
Timeout Error Code execution time exceeds time limit. Runtime Error
Name Error A local or global name is not defined. Runtime Error
Import Error The imported package is not found. Runtime Error
TABLE XIV: Common C Language errors with categories.
Sub-error Description Category
Division by Zero Attempting to divide a number by zero. Runtime Error
TABLE XV: Common Java Language errors with categories.
Sub-error Description Category
Appendix E Supplementary Experiments
E-A Quantitative Evaluation on APPS
E-B Quantitative Evaluation on LeetCode